OpenStudy (anonymous):

what is the equation of the line that is parallel to 8x-5y=2 and goes through the point (-5,-2) A)y=8/5x+6 B) y=8/5x+10 C)y=8/5x-6 D)y=8/5x-10

OpenStudy (anonymous):

that honestly just looks like a bunch of numbers to me

OpenStudy (anonymous):

i dont know how to work these clearly as much as people try to explain them

OpenStudy (shamim):

did u c my work

OpenStudy (anonymous):

i just did right now im trying to figure this out

OpenStudy (shamim):

8x-5y=2 5y=8x-2 y=(8/5)x-2/8 y=(8/5)x-1/4 so slope of ur line is m=8/5

OpenStudy (anonymous):

no im getting a little more yea.

OpenStudy (anonymous):

could you help me with another similar to that one?

OpenStudy (anonymous):

what is the equation of the line that is perpendicular to y=7/9+15 and goes through the point (1,-2) A) 9x+7y=5 B)9x+7y=-5 C)9x+7y=32 D)9x+7y=-23

OpenStudy (anonymous):

yes im sorry

OpenStudy (anonymous):

would it be B?

OpenStudy (anonymous):

y-y1=7/9(x-x1)? is that part right?